A pont charge Q1=5Ng is at the origin and pont charge Q2=3nC is on x-axis at x=3cm. Point P is on the y-axis with coordinate (0,4cm). Calculate the electric fields E1 and E2 at point P due to the charges Q1 and Q2
1
Expert's answer
2020-07-02T17:15:52-0400
"E_1=k\frac{q_1}{r_1^2}\\E_1=9\cdot10^{9}\frac{5\cdot10^{-9}}{0.04^2}=28125\frac{V}{m}"

"E_2=k\frac{q_2}{r_2^2}\\E_2=9\cdot10^{9}\frac{3\cdot10^{-9}}{0.04^2+0.04^2}=8437.5\frac{V}{m}"